OpenStudy (anonymous):

find an equation of the line that contains the given pair of points (23,25),(6,8)

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

Find the slope of a line with these points: y2-y1=m(x2-x1) 8-25=m(6-23) -17=m(-17) m=1 Find the line: y = mx +b 8= (1)(6) + b 8= 6+b b = 2 So the line would be: y=x+2

OpenStudy (anonymous):

first find the slope then use the point slope formula or slope intercept